Suppose a baseball is shot up from the ground straight up with an initial velocity of 32 feet per second. A function can be created by expressing distance above the ground, s, as a function of time, t. This function is s = -16t2 + v0t + s0
• 16 represent 1/2g, the gravitational pull due to gravity (measured in feet per second2).
• v0 is the initial velocity (how hard do you throw the object, measured in feet per second).
• s0 is the initial distance above ground (in feet). If you are standing on the ground, then s0 = 0.
3 Questions:
1)What is the function that describes this problem?
2)The ball will be how high above the ground after 1 second?
3)How long will it take to hit the ground?

Answer by Nate(3500)

1) s = -16t^2 + 32t
2) s = -16(1) + 32(1) = -16 + 32 = 16 feet
3) 0 = -16t^2 + 32t
0 = -t^2 + 2t
0 = t^2 - 2t
0 = t(t - 2)
t = 0 and t = 2
After two seconds, the ball will have hit the ground.